how do you solve the quadratic equation
3x2-7x-6=0?

Answer by sudhanshu_kmr(1152)


3x2-7x-6=0?
=> 3x2-9x + 2x-6=0
=> 3x(x-3)+2(x-3) = 0
=> (3x+2)(x-3) = 0
x = -2/3 or x = 3